# A class of forward-backward regularizations of the Perona-Malik equation with variable exponent

### v2

@@ 3 Weak solutions of the regularized problem (21)–(23)
Then the Euler-Lagrange equation (27) can be equivalently rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\vec{q}(x,\nabla\bar{u}_{m})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (32)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To establish the existence of weak solutions to problem (21)–(23), we use Corollary 1 to obtain uniform estimates for the approximate solutions.
Lemma 12
@@
$\sup_{0\leq t\leq T}\left\|u_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\left\|\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},~\left\|\nabla\sqrt{\varepsilon-\gamma}\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},$ (33)
$\sup_{0\leq t\leq T}\left\|\bar{u}_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\sup_{0\leq t\leq T}\left\|\vec{q}(x,\nabla\bar{u}_{m})\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M_{2},$ (34)
$\left\|u_{m}-\bar{u}_{m}\right\|_{L^{2}(Q_{T})}\leq\frac{1}{m}M_{2}.$ (35)
Proof 4
For any $t\in[0,T)$, there exists a positive integer $j=1,2,\cdots,m$ such that $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$. Given that $\lambda_{m}^{j}(t)\in[0,1)$, it follows from (31) that
@@
and, analogously, we obtain the corresponding estimate for $\bar{u}_{m}$. Furthermore, we derive from (17) and Lemma 1 that
@@
which yields (34). Using (31), we obtain
@@
Hence, we arrive at inequalities (33). To prove (35), it suffices to show that
@@
The proof is thus complete.
By Lemma 2, upon extracting suitable (not relabeled) subsequences, we obtain the following convergences as $m\to\infty$:
$u_{m}\rightharpoonup u\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega))$ (36)
@@
Since $u_{m}-\bar{u}_{m}$ converges strongly to zero in $L^{2}(Q_{T})$, we deduce that $u=\bar{u}$ a.e. in $Q_{T}$. Moreover, $u(t)\in W_{f}^{1,p(\cdot)}(\Omega)$ a.e. in $[0,T]$ follows from (36). Passing to the limit as $m\rightarrow\infty$ in (32), we then obtain
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{\Lambda}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (37)
@@
Lemma 13
Let $m$ denote the index of the subsequence along which the convergences in (36) hold. Then both $\nabla u_{m}$ and $\nabla\bar{u}_{m}$ converge strongly in $L^{2}(0,T;L^{2}(\Omega))$ as $m\to\infty$.
Proof 5
@@
$\leq$ $2\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx+2\left(i-\frac{m}{T}t\right)^{2}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (38)
@@
$=$ $2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt+\frac{2T}{3m}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (39)
For any $t\in[0,T)$, there exists an index $l\in\{1,2,\cdots,m\}$ such that $(l-1)\frac{T}{m}\leq t<l\frac{T}{m}$. For the case $l\geq 2$, from (31), (39), and the Hölder inequality (see Lemma 3), we derive
@@
$\leq$ $2\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{2T^{2}}{3m^{2}\left(\varepsilon-\gamma\right)}C_{1}+\frac{T}{m}C_{1},$ (40)
where $C_{1}>0$ depends only on $\Omega$, $M_{1}$, and $M_{2}$. As for the case $l=1$, the proof details are omitted. Indeed, integrating both sides of (38) with $i=1$ over $(0,t)$ yields a result analogous to (40).
We now analyze the convergence of $\{\nabla u_{m}\}_{m\geq m_{0}}$ and $\{\nabla\bar{u}_{m}\}_{m\geq m_{0}}$ based on the preceding discussion. By choosing the test functions $\zeta=\bar{u}_{m}-u$ in (32) and $u_{m}-u$ in (37), respectively, and then subtracting the resulting identities, we have, for $t\in[0,T]$,
@@
We denote the three terms on the left-hand side of the above equality by $T_{1}$, $T_{2}$, and $T_{3}$, and write $\theta(m)$ for quantities tending to zero as $m\rightarrow\infty$, uniformly with respect to $t$. By (18) and Lemma 6 (e.g., with $\varPsi=\{\chi_{Q_{t}}\vec{q}(x,\nabla u):t\in[0,T]\}$), we obtain
@@
which, together with (40), implies that
@@
$+\varepsilon\int_{0}^{t}\int_{\Omega}\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)dxd\tau.$ (41)
It remains to estimate the last two terms on the right-hand side of (41). On the one hand, for any $t\in[0,T)$, assume that $t\in[(j-1)\frac{T}{m},j\frac{T}{m})$ for some $j=1,2,\cdots,m$. Then
@@
$\leq$ $\frac{TM_{1}\varepsilon}{2m\left(\varepsilon-\gamma\right)}.$ (42)
On the other hand, we choose the test function $\zeta=\xi\eta$ in equation (32), where $\xi\in C_{c}^{1}(\Omega)$ and $\eta\in C^{1}([0,T])$ satisfies $\eta(T)=0$ and $\eta(0)\neq 0$. Using integration by parts, we get
@@
Letting $m\rightarrow\infty$, we have from (36) that
@@
Similarly, substituting the test function $\zeta$ into (37), we obtain
@@
which yields $u(0)=f$ a.e. in $\Omega$. In summary, combining (41) and (42), we deduce
@@
Hence, from (33), (35), and (36), it follows that $T_{3}=\theta(m)$. Taking into account that $T_{1}+T_{2}+T_{3}=0$ and combining this with the above estimates, we obtain
@@
From (40), we further deduce that
$\nabla\bar{u}_{m}\rightarrow\nabla u\text{ strongly in }L^{2}(0,T;L^{2}(\Omega)).$ (43)
The proof of Lemma 13 is complete.
@@
$\vec{\Lambda}(x,t)=\vec{q}\left(x,\nabla u(x,t)\right)\text{ a.e. in }Q_{T}.$ (44)
Proof 6
@@
Considering the global Lipschitz continuity of the mapping $\xi\mapsto\frac{\xi}{1+\left|\xi\right|^{2}}$, we deduce from (43) that
$\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\rightarrow\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\text{ strongly in }L^{2}(Q_{T}).$ (45)
Hence, to prove (44), it suffices to show that
@@
The integral identity (32) can be rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\cdot\nabla\zeta+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)-2}\nabla\bar{u}_{m}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (46)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Letting $m\rightarrow\infty$ in (46) and using (45), we have
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\cdot\nabla\zeta+\delta\vec{\sigma}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0.$ (47)
@@
Furthermore, substituting $\zeta=\bar{u}_{m}$ into (46) yields
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\bar{u}_{m}+\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)}+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}\right)dxdt=0.$ (48)
@@
for a.e. $(x,t)\in Q_{T}$. Combining this with (43), (47), and (48), we deduce that
@@
Thus, applying the standard monotone method, we obtain $\vec{\sigma}=\left|\nabla u\right|^{p(x)-2}\nabla u$ a.e. in $Q_{T}$, which verifies that (44) holds.
Based on the results established earlier, we now prove that the regularized problem (21)–(23) admits a unique weak solution.
Proof of Theorem 2 1
Combining (37) and (44) yields the following integral identity
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{q}(x,\nabla u)\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (49)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. The Neumann boundary condition of problem (21)–(23) is implicitly incorporated in the weak variational formulation $J^{m,j}$ and can be recovered via integration by parts. Next, we address the uniqueness of the weak solution to problem (21)–(23). To this end, assume that $u_{1}$ and $u_{2}$ are two weak solutions with the same initial data $f$. For any $t\in[0,T]$, choosing the test function $\zeta=\left(u_{1}-u_{2}\right)\chi_{[0,t]}(\tau)$ in (49), we obtain
@@
Using Gronwall’s inequality, we derive that $u_{1}=u_{2}$ a.e. in $Q_{T}$, which shows the uniqueness of the weak solution to problem (21)–(23).
To prove Theorem 2, it remains to establish the energy estimate (25). By Lemma 11, for any $\gamma\in(0,\varepsilon)$, there exists $m_{0}(\gamma)>0$ such that, for all $m\geq m_{0}$ and for a.e. $t\in[0,T)$, we derive from (28) the following inequalities, where $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$ holds for some $j$:
@@
From Lemma 13, we know that $\nabla\bar{u}_{m}(t)\rightarrow\nabla u(t)$ a.e. in $\Omega$. Then, Fatou’s Lemma implies that
@@
for a.e. $t\in[0,T]$. This completes the proof of Theorem 2.
4 Proof of Theorem 1
In this section, we prove the existence of Young measure solutions to problem (1)–(3) using Theorem 2 and the corresponding limit process $\varepsilon\to 0^{+}$.
For any $\varepsilon\in(0,1)$, it follows from Theorem 2 that the regularized problem (21)–(23) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, which satisfies
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (50)
@@ 4 Proof of Theorem 1
for a.e. $t\in[0,T]$. Then, from (17), we derive
@@
By Lemmas 1 and 5, along with the mean invariance of $u^{\varepsilon}(t)$, we deduce that
$\left\|u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\left\|\nabla\sqrt{\varepsilon}u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\sup_{0\leq t\leq T}\left\|u^{\varepsilon}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}+\sup_{0\leq t\leq T}\left\|\vec{q}\left(x,\nabla u^{\varepsilon}(t)\right)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M,$ (51)
@@
For a.e. $t\in[0,T]$, let $\nu_{\cdot,t}=(\nu_{x,t})_{x\in\Omega}$ denote the Young measure generated by the sequence $\{\nabla u^{\varepsilon}(t)\}_{0<\varepsilon<1}$. Since the generating sequence satisfies the boundedness condition (14), it follows from Lemmas 8 and 9 that $\left\|\nu_{x,t}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1$ a.e. in $\Omega$. We now verify that the pair $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3). Combining Remark 1 with (51), we infer that
@@
Then, we obtain $\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle$ a.e. in $Q_{T}$, proving (19). Furthermore, using Hölder’s inequality, we derive
@@
Taking the limit $\varepsilon\rightarrow 0^{+}$ in (50), we arrive at
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Moreover, $u(0)=f$ a.e. in $\Omega$ follows from the same argument as in the proof of Lemma 13, and the details are omitted here. Therefore, $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3), and the proof is complete.
Acknowledgements

### v3

@@ 2.3. Definition of Young measure solutions
Then the Euler-Lagrange equation (3.4) can be equivalently rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\vec{q}(x,\nabla\bar{u}_{m})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (3.9)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To establish the existence of weak solutions to problem (2.8)–(2.10), we use Corollary 3.5 to obtain uniform estimates for the approximate solutions.
Lemma 3.6.
@@
$\sup_{0\leq t\leq T}\left\|u_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\left\|\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},~\left\|\nabla\sqrt{\varepsilon-\gamma}\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},$ (3.10)
$\sup_{0\leq t\leq T}\left\|\bar{u}_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\sup_{0\leq t\leq T}\left\|\vec{q}(x,\nabla\bar{u}_{m})\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M_{2},$ (3.11)
$\left\|u_{m}-\bar{u}_{m}\right\|_{L^{2}(Q_{T})}\leq\frac{1}{m}M_{2}.$ (3.12)
Proof.
For any $t\in[0,T)$, there exists a positive integer $j=1,2,\cdots,m$ such that $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$. Given that $\lambda_{m}^{j}(t)\in[0,1)$, it follows from (3.8) that
@@
and, analogously, we obtain the corresponding estimate for $\bar{u}_{m}$. Furthermore, we derive from (2.4) and Lemma 2.3 that
@@
which yields (3.11). Using (3.8), we obtain
@@
Hence, we arrive at inequalities (3.10). To prove (3.12), it suffices to show that
@@
The proof is thus complete. ∎
By Lemma 2.4, upon extracting suitable (not relabeled) subsequences, we obtain the following convergences as $m\to\infty$:
$u_{m}\rightharpoonup u\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega))$ (3.13)
@@
Since $u_{m}-\bar{u}_{m}$ converges strongly to zero in $L^{2}(Q_{T})$, we deduce that $u=\bar{u}$ a.e. in $Q_{T}$. Moreover, $u(t)\in W_{f}^{1,p(\cdot)}(\Omega)$ a.e. in $[0,T]$ follows from (3.13). Passing to the limit as $m\rightarrow\infty$ in (3.9), we then obtain
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{\Lambda}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (3.14)
@@
Lemma 3.7.
Let $m$ denote the index of the subsequence along which the convergences in (3.13) hold. Then both $\nabla u_{m}$ and $\nabla\bar{u}_{m}$ converge strongly in $L^{2}(0,T;L^{2}(\Omega))$ as $m\to\infty$.
Proof.
@@
$\leq$ $2\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx+2\left(i-\frac{m}{T}t\right)^{2}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (3.15)
@@
$=$ $2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt+\frac{2T}{3m}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (3.16)
For any $t\in[0,T)$, there exists an index $l\in\{1,2,\cdots,m\}$ such that $(l-1)\frac{T}{m}\leq t<l\frac{T}{m}$. For the case $l\geq 2$, from (3.8), (3.16), and the Hölder inequality (see Lemma 2.5), we derive
@@
$\leq$ $2\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{2T^{2}}{3m^{2}\left(\varepsilon-\gamma\right)}C_{1}+\frac{T}{m}C_{1},$ (3.17)
where $C_{1}>0$ depends only on $\Omega$, $M_{1}$, and $M_{2}$. As for the case $l=1$, the proof details are omitted. Indeed, integrating both sides of (3.15) with $i=1$ over $(0,t)$ yields a result analogous to (3.17).
We now analyze the convergence of $\{\nabla u_{m}\}_{m\geq m_{0}}$ and $\{\nabla\bar{u}_{m}\}_{m\geq m_{0}}$ based on the preceding discussion. By choosing the test functions $\zeta=\bar{u}_{m}-u$ in (3.9) and $u_{m}-u$ in (3.14), respectively, and then subtracting the resulting identities, we have, for $t\in[0,T]$,
@@
We denote the three terms on the left-hand side of the above equality by $T_{1}$, $T_{2}$, and $T_{3}$, and write $\theta(m)$ for quantities tending to zero as $m\rightarrow\infty$, uniformly with respect to $t$. By (2.5) and Lemma 2.8 (e.g., with $\varPsi=\{\chi_{Q_{t}}\vec{q}(x,\nabla u):t\in[0,T]\}$), we obtain
@@
which, together with (3.17), implies that
@@
$+\varepsilon\int_{0}^{t}\int_{\Omega}\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)dxd\tau.$ (3.18)
It remains to estimate the last two terms on the right-hand side of (3.18). On the one hand, for any $t\in[0,T)$, assume that $t\in[(j-1)\frac{T}{m},j\frac{T}{m})$ for some $j=1,2,\cdots,m$. Then
@@
$\leq$ $\frac{TM_{1}\varepsilon}{2m\left(\varepsilon-\gamma\right)}.$ (3.19)
On the other hand, we choose the test function $\zeta=\xi\eta$ in equation (3.9), where $\xi\in C_{c}^{1}(\Omega)$ and $\eta\in C^{1}([0,T])$ satisfies $\eta(T)=0$ and $\eta(0)\neq 0$. Using integration by parts, we get
@@
Letting $m\rightarrow\infty$, we have from (3.13) that
@@
Similarly, substituting the test function $\zeta$ into (3.14), we obtain
@@
which yields $u(0)=f$ a.e. in $\Omega$. In summary, combining (3.18) and (3.19), we deduce
@@
Hence, from (3.10), (3.12), and (3.13), it follows that $T_{3}=\theta(m)$. Taking into account that $T_{1}+T_{2}+T_{3}=0$ and combining this with the above estimates, we obtain
@@
From (3.17), we further deduce that
$\nabla\bar{u}_{m}\rightarrow\nabla u\text{ strongly in }L^{2}(0,T;L^{2}(\Omega)).$ (3.20)
The proof of Lemma 3.7 is complete. ∎
@@
$\vec{\Lambda}(x,t)=\vec{q}\left(x,\nabla u(x,t)\right)\text{ a.e. in }Q_{T}.$ (3.21)
Proof.
@@
Considering the global Lipschitz continuity of the mapping $\xi\mapsto\frac{\xi}{1+\left|\xi\right|^{2}}$, we deduce from (3.20) that
$\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\rightarrow\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\text{ strongly in }L^{2}(Q_{T}).$ (3.22)
Hence, to prove (3.21), it suffices to show that
@@
The integral identity (3.9) can be rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\cdot\nabla\zeta+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)-2}\nabla\bar{u}_{m}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (3.23)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Letting $m\rightarrow\infty$ in (3.23) and using (3.22), we have
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\cdot\nabla\zeta+\delta\vec{\sigma}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0.$ (3.24)
@@
Furthermore, substituting $\zeta=\bar{u}_{m}$ into (3.23) yields
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\bar{u}_{m}+\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)}+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}\right)dxdt=0.$ (3.25)
@@
for a.e. $(x,t)\in Q_{T}$. Combining this with (3.20), (3.24), and (3.25), we deduce that
@@
Thus, applying the standard monotone method, we obtain $\vec{\sigma}=\left|\nabla u\right|^{p(x)-2}\nabla u$ a.e. in $Q_{T}$, which verifies that (3.21) holds. ∎
Based on the results established earlier, we now prove that the regularized problem (2.8)–(2.10) admits a unique weak solution.
Proof of Theorem 3.1.
Combining (3.14) and (3.21) yields the following integral identity
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{q}(x,\nabla u)\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (3.26)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. The Neumann boundary condition of problem (2.8)–(2.10) is implicitly incorporated in the weak variational formulation $J^{m,j}$ and can be recovered via integration by parts. Next, we address the uniqueness of the weak solution to problem (2.8)–(2.10). To this end, assume that $u_{1}$ and $u_{2}$ are two weak solutions with the same initial data $f$. For any $t\in[0,T]$, choosing the test function $\zeta=\left(u_{1}-u_{2}\right)\chi_{[0,t]}(\tau)$ in (3.26), we obtain
@@
Using Gronwall’s inequality, we derive that $u_{1}=u_{2}$ a.e. in $Q_{T}$, which shows the uniqueness of the weak solution to problem (2.8)–(2.10).
To prove Theorem 3.1, it remains to establish the energy estimate (3.2). By Lemma 3.4, for any $\gamma\in(0,\varepsilon)$, there exists $m_{0}(\gamma)>0$ such that, for all $m\geq m_{0}$ and for a.e. $t\in[0,T)$, we derive from (3.5) the following inequalities, where $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$ holds for some $j$:
@@
From Lemma 3.7, we know that $\nabla\bar{u}_{m}(t)\rightarrow\nabla u(t)$ a.e. in $\Omega$. Then, Fatou’s Lemma implies that
@@
for a.e. $t\in[0,T]$. This completes the proof of Theorem 3.1. ∎
4. Proof of Theorem 1.1
In this section, we prove the existence of Young measure solutions to problem (1.1)–(1.3) using Theorem 3.1 and the corresponding limit process $\varepsilon\to 0^{+}$.
For any $\varepsilon\in(0,1)$, it follows from Theorem 3.1 that the regularized problem (2.8)–(2.10) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, which satisfies
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (4.1)
@@
for a.e. $t\in[0,T]$. Then, from (2.4), we derive
@@
By Lemmas 2.3 and 2.7, along with the mean invariance of $u^{\varepsilon}(t)$, we deduce that
$\left\|u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\left\|\nabla\sqrt{\varepsilon}u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\sup_{0\leq t\leq T}\left\|u^{\varepsilon}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}+\sup_{0\leq t\leq T}\left\|\vec{q}\left(x,\nabla u^{\varepsilon}(t)\right)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M,$ (4.2)
@@
For a.e. $t\in[0,T]$, let $\nu_{\cdot,t}=(\nu_{x,t})_{x\in\Omega}$ denote the Young measure generated by the sequence $\{\nabla u^{\varepsilon}(t)\}_{0<\varepsilon<1}$. Since the generating sequence satisfies the boundedness condition (2.1), it follows from Lemmas 2.10 and 2.12 that $\left\|\nu_{x,t}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1$ a.e. in $\Omega$. We now verify that the pair $\left(u,\nu\right)$ is a Young measure solution to problem (1.1)–(1.3). Combining Remark 2.13 with (4.2), we infer that
@@
Then, we obtain $\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle$ a.e. in $Q_{T}$, proving (2.6). Furthermore, using Hölder’s inequality, we derive
@@
Taking the limit $\varepsilon\rightarrow 0^{+}$ in (4.1), we arrive at
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Moreover, $u(0)=f$ a.e. in $\Omega$ follows from the same argument as in the proof of Lemma 3.7, and the details are omitted here. Therefore, $\left(u,\nu\right)$ is a Young measure solution to problem (1.1)–(1.3), and the proof is complete.
Acknowledgment